# Building Trustworthy AI by Addressing its 16+2 Desiderata with Goal-Directed Commonsense Reasoning

### v1

@@ 3 Addressing 16+2 desiderata with s(CASP)
The term abduction refers to a form of reasoning that is concerned with the generation and evaluation of explanatory hypotheses. We could also think of abduction as assumptions-based reasoning. Abductive reasoning leads back from facts to a proposed explanation of those facts or assumptions that will explain that fact.
s(CASP)
@@
More formally, abduction is a form of reasoning where, given the premise P $\Rightarrow$\$ Q, and the observation Q, one surmises (abduces) that P holds. More generally, given a theory T, an observation O, and a set of abducibles A, then E is an explanation of O (where E $\subset$\$ A) if:
\$T $\cup$ E $\models$ O\$
\$T $\cup$ E\$ is consistent
We can think of abducibles A as a set of assumptions. Generally, A consists of a set of propositions such that if p $\in$\$ A, then there is no rule in theory T with p as its head (that is, there is no way to argue for p).
@@
We assume the theory T to be an answer set program. Under a goal-directed execution regime, an ASP system can be extended with abduction by simply adding the following rule for an abducible p:
p :- not neg_p.
neg_p :- not p.
@@
Two of the desiderata deal purely with the way an AI system communicates with a user (this and LABEL:des:15). Theory of Mind involves building and continually updating nuanced models of conversation partners, including their knowledge, intentions, and communication styles, to interact appropriately. This helps the AI decide how terse or elaborate to be, when to ask clarifying questions, and how to adapt to ambiguity while avoiding miscommunication. Additionally, the AI must maintain a self-model to understand its capabilities, limitations, and role in the interaction.
s(CASP)
@@
While Lenat and Marcus posed the combination of LLMs and Cyc as a potential solution for the shortcomings of both, that was purely an idea for the future. However, the neuro-symbolic approach of combining LLMs with s(CASP) has already been explored for a variety of domains (??). In this use case, the LLM provides a natural language semantic parser for a s(CASP) reasoner. This constrains the LLM to increase response truthfulness and fluency, and allows for easy user-interfacing with a symbolic system. Unlike other methods to improve LLM accuracy (such as chain-of-thought prompting (?) or more intensive training (?)), using a s(CASP) reasoner provides actual, logically-grounded constraints. The s(CASP) system also determines the boundary of the “cognition” of AI with a dynamically maintained knowledge base, which contributes as an essential part to a self-model, i.e., to avoid hallucination and be aware that something beyond its capability exists. When the current dialogue is far beyond the task at hand (e.g., in concierge chatbot for restaurant recommendation (?)), it will be labeled as “irrelevant” by the LLM, driving the conversation back to the task. Additionally, attaching an LLM to s(CASP) allows s(CASP) to tailor its language according to the theory of mind as well as an LLM can.
@@
Discussion: LLMs are very adaptable to the user, having been trained on numerous styles of text and speaking. Research finds that LLMs can automatically moderate their responses based on the tone (both spoken and written) of the user’s input (??). However, unconstrained LLMs by themselves can be easily “jailbroken” to speak in inappropriate ways for the audience (?).
[…]
Quantifier-Fluency is the ability to understand ambiguous data, or data where information is missing. In logical systems, unknown variables are understood by being “quantified”, that is, for every variable that is not known we assign it a value that is known. In answer set programs, this typically takes the form of grounding. However, not every variable can be ground and not everything can be known. Truly trustworthy AI would need to be able to reason with incomplete state.
s(CASP)
Using abducibles in goal-directed ASP, information that is unknown can be guessed (possibly creating multiple worlds) in order to find a solution to a given problem. For example, the following pattern models whether the documents we have to certify that we are a large_family are valid or not:
large_family :- evidence(large_family).
neg_large_family:- evidence(-large_family).
large_family :-
not evidence(large_family), not neg_large_family.
neg_large_family :-
not evidence(-large_family), not large_family.
So we avoid introducing that information, and the reasoner would reason assuming both scenarios. Lines 1 and 2 capture the case where we have evidence for or against, and lines 3-6 create a loop otherwise so that we assume one option or the other during evaluation.
This has been exercised with ambiguities in the legal field by ? (?) and ? (?).
@@
Discussion: The way LLMs handle quantifier-fluency is by leaning on prior knowledge to make educated guesses. However, it can wrongly fill in unknown information with unrelated assumptions. Cyc also relies on the completeness of its knowledge base for reasoning, and thus cannot reason without information that is not in or deducible/inducible from its knowledge base.
[…]
@@
Modal-fluency is the ability to understand and use qualified statements expressing things that may or may not be ultimately true. This includes statements such as “He hopes that it was successful” or “I believe she’s afraid that it may cost her the job”, which can be difficult for an AI system to reason over to determine whether something did or did not actually happen.
s(CASP)
@@
ASP is based on stable model semantics that allow for multiple possible worlds. The semantics of modal logic is also given in terms of possible worlds. Thus, modal logics can be elegantly modeled with ASP/s(CASP). The $\neg K$ operator in epistemic modal logic, e.g., is nothing but default negation. Note that the concepts of “p is definitely true”, “p may be true”, “p is unknown”, “p may be false”, “p may be unknown”, “p is definitely false” are represented by ASP expressions“p”, “not -p”, “not -p $\wedge$\$ not p”, “not p”, and “-p”, respectively, where -p represents strong negation (?). Deontic logic, or the modal logic of obligations, can also be naturally represented in ASP/s(CASP). An obligation (the OB modal operator) can be represented as ASP’s global constraint. In fact, since the odd loop over negation represents a global constraint, it can be used to elegantly resolve the contrary-to-duty paradox (as well as other paradoxes) of deontic logic. Consider the following statements for a housing community: (1) It ought to be the case that there are no dogs. (2) It ought to be the case that if there are no dogs, then there are no warning signs. (3) If there are dogs, then it ought to be the case that there are warning signs. (4) There are dogs. This can be represented as:
(1) dog :- not -dog, not dog.
(2) :- -dog, sign.
(3) sign :- dog. :- dog, not sign.
(4) dog.
@@
Given this program, no world will be permitted in which there isn’t a beware-of-dog sign. If we remove (4), then no worlds will be possible in which a dog is present, as now constraint (1) will kick in. This ASP encoding thus elegantly resolves the contrary-to-duty paradox (a.k.a. the Chisholm paradox).
@@
Discussion: Although LLMs do not possess or store any knowledge separately, the amount of natural language they’ve internalized gives them some ability to “understand” complex sentences. Larger LLMs perform moderately well with modal logic, though there is still much room for improvement (?). While the LLMs can parse modal sentences, they lack a true understanding of the logic behind modal statements. As with analogical reasoning (Desiderata LABEL:des:04), replicating response patterns to modal logic does not equate to an understanding of it, which makes them prone to strange mistakes and hallucinations. This can be contrasted with the fully symbolic representation of Cyc, where modal logic is represented in its hand-crafted axioms (if something is “believed” it means that the “believer” thinks it is true, for example).
[…]
@@
Defeasibility represents the ability to acquire new information and change previously held beliefs, many of which were likely only true by default in the first place. This requires the ability to reconsider previous beliefs and evaluate them on their own merits against new information before deciding whether the new or old information should be kept.
s(CASP)
Default rules can be modified by induction with exceptions, further refining the criteria for deduction. FOLD primarily learns in the form of default rules and exceptions (as in Desiderata LABEL:des:03). By using this format, s(CASP) is able to incorporate the defeasibility desiderata. As new data is integrated, information can be marked as an exception, or even an exception to an exception, allowing for a high level of fidelity even when receiving new information. Because the form the knowledge takes is default rules in a logic program, it is easy to see what was learned and correct it if there are obvious mistakes. Thus, when a constraint cannot be learned, it can be added to guarantee correct behavior.
@@
In the reliable chatbot mentioned in desiderata LABEL:des:06, s(CASP) maintains the state of the conversation, tracking the progress of the task. Once new information comes, s(CASP) first checks if the is is consistent with the current state and knowledge base, and then updates it to the state. Inconsistent information will either be discarded or returned to the conversation partner for further clarification.
@@
Discussion: Both LLMs and Cyc have a similar approach to defeasibility, which is to yield to human input whenever given. For Cyc, this is a part of the continual hand-grooming of its large knowledge base. When information is added that conflicts with existing information, a human de-conflicts the two either by removing one piece of information or adding additional context around it. LLMs have a less permanent approach, wherein rather than removing data from its memory (a noted difficulty with LLMs (?)) the LLM generates new responses when challenged. While human input is almost always correct relative to AI output, both systems do lack the ability to critically evaluate their answers themselves (?).
[…]
@@
Another facet of XAI represented in Lenat’s desiderata is the “pro and con arguments” desiderata, which is a succinct description of the ability to explain why a decision may or may not be true. This covers the ambiguity of life, where things may not be objectively true but rather situationally true based on some confounding factors. Additionally, in any given situation there may be multiple answers that are correct enough.
s(CASP)
@@
Answer set programs inherently generate pros and cons in the form of multiple worlds. When looking at a solution in one possible world, s(CASP) outlines which factors must be true or false to support that world. Additionally, if a proof of a goal g fails, one can immediately query not g (which must succeed in s(CASP)) to understand the cause of failure (?). Similarly, counterfactuals can be computed in s(CASP) with ease (?) to examine various what-if scenarios.
@@
Discussion: This functionality is built into Cyc which provides arguments for and against each point it provides, along with heuristics for those arguments, at increasingly granular levels. LLMs, however, do not provide balanced pro and con arguments because they fold too easily against user disagreement (see Desiderata LABEL:des:09). Being unable to defend a premise when challenged makes it impossible to trust a fair evaluation of its own answers (?).
[…]
@@
Contextual understanding allows an AI to interpret and adapt knowledge, behavior, and communication based on situational, cultural, and temporal factors, such as knowing when cheering is appropriate or recognizing shifting truths over time. It must reason both within a context (e.g., belief systems or cultural norms) and about contexts (e.g., deciding whether an inference holds true in multiple settings).
s(CASP)
@@
The multiple worlds reasoning of all ASP programs allows for the exploration of multiple contexts. Consider this example from Peter Norvig:
1 People can talk.
2 Non-human animals are not able to talk.
3 Human-like cartoon characters can talk.
4 Fish can swim.
5 A fish is a non-human animal.
6 Nemo is a human-like cartoon character.
7 Nemo is a fish.
@@
The solution to this problem is difficult to represent in classical logic, but simple in ASP. In ASP there would be two mutually exclusive worlds, a “cartoon” world where Nemo can talk, and a “real” world where Nemo cannot. Of course, context can also be carried as an argument of a predicate in s(CASP).
@@
Discussion: Cyc contains over 10,000 contexts that are nested within other contexts to delineate topics, time periods, real or not real, and more. This allows for a robust differentiation of contexts for answers that fit the situation. However, it is still susceptible to encountering contexts not present in its handmade database and slowdowns related to having to go down multiple different context chains to provide answers. Still, this is better than LLMs which have difficulty keeping context even from a single conversation if the conversation grows too large (?).
[…]
@@
Meta-knowledge and meta-reasoning enable an AI to reflect on its own knowledge, assess the reliability and source of its beliefs, and understand its strengths and limitations in performing tasks. This includes the ability to introspect, to adapt strategies mid-process, to account for changes in reasoning over time, and to critically evaluate the reliability of its sources.
s(CASP)
@@
Negations as failure gives us a form of meta reasoning because it allows us to reason about and using a failed proof, enabling correction. The co-inductive hypothesis set (CHS) maintained in s(CASP) (?) keeps track of everything that has been found to be true, further allowing it to solve queries about the proof process itself. Did we solve this through assumption (co-inductive) or was the reasoning well-founded (inductive)? Meta-programming and meta-reasoning can also be explicitly incorporated in s(CASP) by making the CHS accessible to the knowledge engineer through appropriate built-ins.
@@
Discussion: This function is available in a limited capacity in both LLMs and Cyc. LLMs can be asked whether their answer is correct or not, but is prone to over-correcting at any user feedback. Cyc has reasoners about its reasoning, but because its knowledge base is human-constructed it takes the contents of its knowledge base as facts.
[…]
@@
A trustworthy AI must adhere to a transparent ethical framework, guided by strong core principles. Crucially, the AI’s ethical commitments should be visible and understandable to those it affects, forming an unchanging part of its foundational contract with users.
s(CASP)
It is easy for users to add constraints to a s(CASP) program, forcing it to be explicitly ethical. This is unlike deep learning systems which are difficult to correct or add constraints that are not in the data they were trained on.
? (?) have used s(CASP) constraints to validate a patient controlled analgesia (PCA) pump modeled with event calculus, and ? (?) in the automation of administrative processes. These two applications illustrate how ethical restrictions can be modeled as constraints.
@@
Discussion: As LLMs are still often “jailbroken” to provide data against the terms of their use, it is clear that this principle is failed by LLMs. As for Cyc, without access to their systems it is hard to verify if it is explicitly ethical or not. However, therein lies a lack of transparency that does not at first glance allow that foundational ethical contract to be formed with it.
[…]
@@
An effective AI system must respond to a user query in a reasonable length of time. While in rare cases a long processing time may be acceptable, in most use cases users expect answers to their queries in a matter of seconds.
s(CASP)
@@
Rather than spending a large time training, as with deep learning approaches, logic programming systems like s(CASP) perform their reasoning when the query is executed. Reasoning in real world environments with many objects, each having state and properties, becomes overwhelming and can result in long execution time for queries. To avoid this problem, many logic programs are kept small and domain specific. Thus these programs lack important context for reasoning.
@@
The Virtually Embodied Common Sense Reasoning (VECSR) system is built on a foundation of s(CASP) reasoning to solve problems in simulated embodied environments (?). The advantage of connecting a s(CASP)-based solving system to a virtual simulation (as with the other connections to LLMs and data mentioned above) is that there is ultimate access to the context required for decision making. However, this results in a large knowledge base and long durations for query-execution. The VECSR system employs several compile-time static analysis tools that can bring the reasoning time in even large programs to under a second much of the time, faster than even LLMs. And unlike LLMs, logical constraints can be used to guarantee correctness and executability that LLMs can’t provide (?). Using a unified and multi-domain reasoner like s(CASP) also eliminates the need for switching between multiple reasoners. Thus, s(CASP) can be used for real-world, real-time reasoning with a high degree of accuracy.
@@
Sufficiently Lingual and Embodied refers to the ability of an AI system to communicate sufficiently for the task at hand. And while in some applications, the minimum necessary linguistics may not include natural language at all (such as in systems designed to produce numerical answers for math problems), in most use-cases some natural language is optimal for human understanding. Even when answers can be provided with little language, it is important to be able to explain “why” an answer is the case. In most cases, knowledge and reasoning are language-independent, but a trustworthy AI will nonetheless be able to communicate well.
s(CASP)

### v2

@@ 1 Multiple Possible Worlds ! ASP Multiple Worlds[6]
#05 Abductive Reasoning: The term abduction refers to a form of reasoning that is concerned with the generation and evaluation of explanatory hypotheses. We could also think of abduction as assumptions-based reasoning. Abductive reasoning leads back from facts to a proposed explanation of those facts or assumptions that will explain that fact.
@@
s(CASP): More formally, abduction is a form of reasoning where, given the premise P $\Rightarrow$$ Q, and the observation Q, one surmises (abduces) that P holds. More generally, given a theory T, an observation O, and a set of abducibles A, then E is an explanation of O (where E $\subset$$ A) if:
\$T $\cup$ E $\models$ O\$ \$T $\cup$ E\$ is consistent
@@
We can think of abducibles A as a set of assumptions. Generally, A consists of a set of propositions such that if p $\in$$ A, then there is no rule in theory T with p as its head (that is, there is no way to argue for p). We assume the theory T to be an answer set program. Under a goal-directed execution regime, an ASP system can be extended with abduction by simply adding the following rules for an abducible p:
p :- not neg_p. neg_p :- not p.
@@
#06 Theory of Mind: Two of the desiderata deal purely with the way an AI system communicates with a user (this and 1). Theory of Mind involves building and continually updating nuanced models of conversation partners, including their knowledge, intentions, and communication styles, to interact appropriately. This helps the AI decide how terse or elaborate to be, when to ask clarifying questions, and how to adapt to ambiguity while avoiding miscommunication. Additionally, the AI must maintain a self-model to understand its capabilities, limitations, and role in the interaction.
@@
s(CASP): While Lenat and Marcus posed the combination of LLMs and Cyc as a potential solution for the shortcomings of both, that was purely an idea for the future. However, the neuro-symbolic approach of combining LLMs with s(CASP) has already been explored for a variety of domains (Zeng et al., 2023; Basu et al., 2021a). In this use case, the LLM provides a natural language semantic parser for a s(CASP) reasoner. This constrains the LLM to increase response truthfulness and fluency, and allows for easy interfacing with a symbolic system. Unlike other methods to improve LLM accuracy (such as chain-of-thought prompting or more intensive training), using a s(CASP) reasoner provides actual logically-grounded constraints. The s(CASP) system also determines the boundary of the “cognition” of AI with a dynamically maintained knowledge base, which contributes as an essential part to a self-model, i.e., to avoid hallucination and be aware that something beyond its capability exists. When the current dialogue is far beyond the task at hand it will be labeled as “irrelevant” by the LLM-based semantic parser, driving the conversation back to the task. Additionally, attaching an LLM to s(CASP) allows s(CASP) to tailor its language according to the theory of mind as well as an LLM can.
@@
Discussion: LLMs are very adaptable to the user, having been trained on numerous styles of text and speaking. Research finds that LLMs can automatically moderate their responses based on the tone (both spoken and written) of the user’s input (Lin et al., 2024). However, unconstrained LLMs by themselves can be easily “jailbroken” to speak in inappropriate ways (Chao et al., 2023).
#07 Quantifier-fluency: Quantifier fluency relates to being able to reason correctly with quantifiers. Quantifiers play a crucial role in interpreting meaning of sentences. For example, as Lenat & Marcus (2023) point out, the sentence “Every Swede has a king” versus the sentence “Every Swede has a mother” have to be interpreted appropriately.
s(CASP): Since s(CASP) supports predicates, quantifiers can be easily modeled. In an answer set program, all variable are universally quantified. However, negation as failure can introduce universally quantified variables in the body of the rules. Consider for example, the following definition of a bachelor.
bachelor(X) :- man(X), not married_status(X).
married_status(X) :- married(X,Y,T). % X married Y at time T
The variables Y and T are existentially quantified in the body of the second rule. The negation of married_status in the first rule leads to these variables being universally quantified in the body. That is, X is a bachelor if X is a man, and did not marry anyone at anytime. Thus, goals with universal quantification can appear in the body of s(CASP) programs, allowing it to incorporate both types of quantifiers.
With respect to the example where we have to distinguish between every Swede having a king vs having a mother, such nuances can be modeled via ASP’s global constraints, as illustrated below.
false :- swede(X), not hasking(X). %Every Swede has a king (or, no Swede does not have a king)
hasking(X) :- king(Y, X).
false :- swede(X), swede(Y), X $\neq$$ Y, king(K1, X), king(K2, Y), K1 $\neq$$ K2. % All Swedes have same king
false :- king_of_sweden(K1), king_of_sweden(K2), K1 $\neq$$ K2. % There is only one King of Sweden
king(K,X) :- king_of_sweden(K), swede(X).
false :- swede(X), not hasmother(X). % Every Swede has a mother (or no Swede does not have a mother)
hasmother(X) :- mother(Y, X).
@@
Discussion: The way LLMs handle quantifier fluency is by leaning on prior knowledge to make educated guesses. However, it can wrongly fill in unknown information with unrelated assumptions. Cyc also relies on the completeness of its knowledge base for reasoning, and thus cannot reason without information that is not in or deducible/inducible from its knowledge base.
@@
#08 Modal-fluency: Modal-fluency is the ability to understand and use qualified statements. This includes statements such as “He hopes that it was successful” or “I believe she’s afraid that it may cost her the job”, which can be difficult for an AI system to reason over.
@@
s(CASP): ASP is based on stable model semantics that allow for multiple possible worlds. The semantics of modal logics that can be used to model modal-fluency is also given in terms of possible worlds. Thus, modal logics can be elegantly modeled with ASP/s(CASP). The $\neg K$ operator in epistemic modal logic, e.g., is nothing but default negation. Note that the concepts of “p is definitely true”, “p may be true”, “p is unknown”, “p may be false”, “p may be unknown”, “p is definitely false” are represented by ASP expressions“p”, “not -p”, “not -p $\wedge$$ not p”, “not p”, and “-p”, respectively, where -p represents strong negation (Gelfond & Kahl, 2014). Deontic logic, or the modal logic of obligations, can also be naturally represented in ASP/s(CASP). An obligation (the OB modal operator) can be represented as ASP’s global constraint. In fact, since the odd loop over negation represents a global constraint, it can be used to elegantly resolve the contrary-to-duty paradox (as well as other paradoxes) of deontic logic. Consider the following statements for a housing community:
  1. 1.
    It ought to be the case that there are no dogs.
  2. 2.
    It ought to be the case that if there are no dogs, then there are no warning signs.
  3. 3.
    If there are dogs, then it ought to be the case that there are warning signs.
  4. 4.
    There are dogs.
In our framework, this is simply modeled as:
s(CASP): dog
dog :- not -dog, not dog.
:- -dog, not -warning_sign.
:- dog, not warning_sign.
dog.
@@
Given this program, no world will be permitted in which there is no warning sign. If we remove (4), then no worlds will be possible in which a dog is present, as now constraint (1) will kick in. This ASP encoding thus elegantly resolves the contrary-to-duty paradox (a.k.a. the Chisholm paradox). For a more detailed discussion of the way s(CASP) allows for deontic logic representation, see Gupta et al. (2025)
@@
Discussion: Although LLMs do not possess or store any knowledge separately, the amount of natural language they’ve internalized gives them some ability to “understand” complex sentences. Larger LLMs perform moderately well with modal logic, though there is still much room for improvement (Holliday et al., 2024). While the LLMs can parse modal sentences, they lack a true understanding of the logic behind modal statements. As with analogical reasoning (Desideratum 1), replicating response patterns to modal logic does not equate to an understanding of it, which makes them prone to strange mistakes and hallucinations.
@@
#09 Defeasibility: Defeasibility represents the ability to acquire new information and change previously held beliefs, many of which were likely only true by default in the first place. This requires the ability to reconsider previous beliefs and evaluate them on their own merits against new information before deciding whether the new or old information should be kept.
s(CASP): Defeasibility can be implemented using default rules. Default rules are directly supported in ASP and s(CASP) with the help of negation-as-failure. Also, as discussed earlier, the FOLD algorithm primarily learns a model from data in the form of default rules with exceptions (as discussed in Desideratum 1). These learned rules not only include exception, they may also include exception to an exception, exception an exception to an exception, and so on. This allows for a high level of fidelity when learning or representing a nuanced concept.
Note that default rules also allow ASP and s(CASP) to model human judgment. Consider a doctor prescribing a medicine M for disease D to patient P. Medicine M may have a side-effect for patient P, and so the doctor has to make a judgment call before prescribing. The doctor could aggressively jump to the default conclusion and may just ignore the possibility of a side-effect (contraindication).
prescribe(M, D, P) :- cures(M, D), not contraindicated(M, P).
contraindicated(M, P) :- has_side_effects(M, P).
Or, the doctor could be conservative in jumping to the default:
prescribe(M, D, P) :- cures(M, D), not contraindicated(M, P).
contraindicated(M, P) :- not -has_side_effects(M, P).
The rules above could be simplified and written, respectively, as:
prescribe(M, D, P) :- cures(M, D), not has_side_effects(M, P). %aggressive
prescribe(M, D, P) :- cures(M, D), -has_side_effects(M, P). %conservative
The first rule states that if there is no information about a possible side-effect, then proceed with prescribing. The second rule states that the possibility of a side-effect must be ruled out before prescribing. The ability to represent these nuanced judgment calls using default and strong negation is a positive aspect of ASP and s(CASP).
@@
In the reliable chatbot mentioned in Desideratum 1, s(CASP) maintains the state of the conversation, tracking the progress of the task. Once new information comes, s(CASP) first checks if it is consistent with the current state and knowledge base, and then uses it to update the state. Inconsistent information will either be discarded or clarification will be sought from the human user.
@@
Discussion: Both LLMs and Cyc have a similar approach to defeasibility, which is to yield to human input whenever given. For Cyc, this is a part of the continual hand-grooming of its large knowledge base. When information is added that conflicts with existing information, a human de-conflicts the two either by removing one piece of information or adding additional context around it. Cyc also supports default rules. LLMs have a less permanent approach, wherein rather than removing data from its memory (a noted difficulty with LLMs (Blanco-Justicia et al., 2025)), the LLMs generates new responses when challenged. While human input is almost always correct relative to AI output, both systems do lack the ability to critically evaluate their answers themselves (Antoniou & Batsakis, 2023).
@@
#10 Pro and Con Arguments: Another facet of XAI represented in Lenat’s desiderata is the “pro and con arguments” desideratum, which is a succinct description of the ability to explain why a decision may or may not be true. This covers the ambiguity of life, where things may not be objectively true but rather situationally true based on some confounding factors. Additionally, in any given situation there may be multiple answers that are correct enough.
@@
s(CASP): Answer set programs inherently generate pros and cons in the form of multiple worlds. When looking at a solution in one possible world, s(CASP) outlines which factors must be true or false to support that world. Additionally, if a proof of a goal g fails, one can immediately query not g (which must succeed in s(CASP)) to understand the cause of failure (Murugesan et al., 2024). Similarly, counterfactuals can be computed in s(CASP) with ease (Dasgupta et al., 2025) to examine various what-if scenarios.
@@
Discussion: This functionality is built into Cyc which provides arguments for and against each point it considers, along with heuristics for those arguments, at increasingly granular levels. LLMs, however, do not provide balanced pro and con arguments because they fold too easily against user disagreement (see Desideratum 1). Being unable to defend a premise when challenged makes it impossible to trust a fair evaluation of its own answers (Wang et al., 2023).
@@
#11 Contexts: Contextual understanding allows an AI to interpret and adapt knowledge, behavior, and communication based on situational, cultural, and temporal factors, such as knowing when cheering is appropriate or recognizing shifting truths over time. It must reason both within a context (e.g., belief systems or cultural norms) and about contexts (e.g., deciding whether an inference holds true in multiple settings).
@@
s(CASP): The multiple worlds reasoning of all ASP programs allows for the exploration of multiple contexts. Consider this example from Peter Norvig:
  1. 1.
    People can talk.
  2. 2.
    Non-human animals are not able to talk.
  3. 3.
    Human-like cartoon characters can talk.
  4. 4.
    Fish can swim.
  5. 5.
    A fish is a non-human animal.
  6. 6.
    Nemo is a human-like cartoon character.
  7. 7.
    Nemo is a fish.
@@
The solution to the problem of whether Nemo can talk is difficult to represent in classical logic, but simple in ASP. In ASP there would be two mutually exclusive worlds, a “cartoon” world where Nemo can talk, and a “real” world where Nemo cannot. Of course, context can also be carried as an argument of a predicate in s(CASP).
@@
Discussion: Cyc contains over 10,000 contexts that are nested within other contexts to delineate topics, time periods, real or not real, and more. This allows for a robust differentiation of contexts for answers that fit the situation. However, it is still susceptible to encountering contexts not present in its handmade knowledge base and slowdowns related to having to go down multiple different context chains to provide answers. Still, this is better than LLMs which have difficulty keeping context even from a single conversation if the conversation grows too large (Hatalis et al., 2024).
@@
#12 Meta-Knowledge and Meta-Reasoning: Meta-knowledge and meta-reasoning enable an AI to reflect on its own knowledge, assess the reliability and source of its beliefs, and understand its strengths and limitations in performing tasks. This includes the ability to introspect, to adapt strategies mid-process, to account for changes in reasoning over time, and to critically evaluate the reliability of its sources.
@@
s(CASP): Negation-as-failure gives us a form of meta reasoning because it allows us to reason about a failed proof, enabling correction. The co-inductive hypothesis set (CHS) maintained in s(CASP) (Arias et al., 2018) keeps track of everything that has been found to be true, further allowing it to solve queries about the proof process itself. Did we solve this through assumption (co-inductive) or was the reasoning well-founded (inductive)? Meta-programming and meta-reasoning can also be explicitly incorporated in s(CASP) by making the CHS accessible to the knowledge engineer through appropriate built-ins.
@@
Discussion: This function is available in a limited capacity in both LLMs and Cyc. LLMs can be asked whether their answer is correct or not, but are prone to over-correcting if any user feedback is given. Cyc has reasoners about its reasoning, and thus can perform meta-reasoning.
@@
#13 Explicitly Ethical: Trustworthy AI must adhere to a transparent ethical framework, guided by strong core principles. Crucially, the AI’s ethical commitments should be visible and understandable to those it affects, forming an unchanging part of its foundational contract with users.
s(CASP): It is possible for users to add constraints to a s(CASP) program, forcing it to be explicitly ethical. This is unlike deep learning systems which are difficult to correct or add constraints that are not in the data they were trained on. Vašíček et al. (2024) have used s(CASP) constraints to validate a patient controlled analgesia (PCA) pump modeled with event calculus, and Arias et al. (2024b) used s(CASP) in the automation of administrative processes. These two applications illustrate how ethical restrictions can be modeled as constraints. Additionally, recently Gupta et al. (2025) demonstrated that there is 1-1 mapping between modal logics and ASP, and how norms of obligation, impermissibility, and permission of deontic modal logic can be modeled elegantly in s(CASP) through odd loop over negation. Conditional norms as well as conflicting norms can be modeled. This expressiveness is crucial for representing ethical systems. Additionally, s(CASP) is fully open source, allowing for complete review of its source code.
@@
Discussion: As LLMs are still often “jailbroken” to provide data against the terms of their use, it is clear that this principle is failed by LLMs. As for Cyc, without access to their systems it is hard to verify if it is explicitly ethical or not.
@@
#14 Sufficient Speed: An effective AI system must respond to a user query in a reasonable length of time. While in rare cases a long processing time may be acceptable, in most use cases users expect answers to their queries in a matter of seconds.
@@
s(CASP): Rather than spending a large time training, as with deep learning approaches, logic programming systems like s(CASP) perform their reasoning when the query is executed. Reasoning in real world environments with many objects, each having state and properties, becomes overwhelming and can result in long execution time for queries. To avoid this problem, many logic programs are kept small and domain specific. Thus these programs lack important context for reasoning.
@@
The Virtually Embodied Common Sense Reasoning (VECSR) system is built on a foundation of s(CASP) reasoning to solve problems in simulated embodied environments (Tudor et al., 2025). The advantage of connecting a s(CASP)-based system to a virtual simulation (as with the other connections to LLMs and data mentioned above) is that there is ultimate access to the context required for decision making. However, this results in a large knowledge base and long durations for query-execution. The VECSR system employs several compile-time static analysis techniques that can bring the reasoning time in even large programs to under a second much of the time, faster than even LLMs (proven in Tudor et al. (2025)). And unlike LLMs, logical constraints can be used to guarantee correctness and executability. Using a unified and multi-domain reasoner like s(CASP) also eliminates the need for switching between multiple reasoners. Thus, s(CASP) can be used for real-world, real-time reasoning with a high degree of accuracy.
@@
#15 Sufficiently Lingual and Embodied: Sufficiently Lingual and Embodied refers to the ability of an AI system to communicate sufficiently for the task at hand. And while in some applications, the minimum necessary linguistics may not include natural language at all (such as in systems designed to produce numerical answers for math problems), in most use-cases some natural language is optimal for human understanding. Even when answers can be provided with little language support, it is important to be able to explain “why” an answer was given. In most cases, knowledge and reasoning are language-independent, but a trustworthy AI will nonetheless be able to communicate well.